# USTC-ICTS/PCFT-24-43

### v2

@@ Appendix A Overlap of a generic pair of $X^{a}_{I}$ and $X^{b}_{I^{\prime}}$
For each pair $a\neq b$, we can first redefine $\psi_{i}$ appropriately such that the relative overlapping is equivalent to $X_{I}^{1}$ and $X_{I^{\prime}}^{2}$. To see this, consider $X_{I}^{a}$ and $X_{I^{\prime}}^{a+b}$ with $b>0$. We can first redefine $\psi_{2i}\rightarrow\psi_{2i-2(a-1)}$ to shift $X_{I}^{a}\rightarrow X_{I}^{1}$ and $X_{I}^{a+b}\rightarrow X_{I}^{1+b}$. We first assume $N$ is coprime with any number between $2$ and $d-1$. Then, we redefine $\psi_{2i-1}\rightarrow\psi_{2bi-(2b-1)}$ and $\psi_{2i}\rightarrow\psi_{2bi-(2b-2)}$. Since $b$ takes values between $1$ and $d-1$, by our coprime assumption this redefinition is a permutation with one cycle. Since random overlapping does not care about the explicit numbering of Majorana fermions, this problem is equivalent to the overlapping between $X_{I}^{1}$ and $X_{I^{\prime}}^{2}$.
For a generic case where $N$ is not coprime with a number between $2$ and $d-1$. The second redefinition does not lead to the overlapping problem between $X_{I}^{1}$ and $X_{I^{\prime}}^{2}$ but $b$ copies of overlapping problem with $N\rightarrow N/b$. To be precise, assume $N/b\in\mathbb{Z}$ and there exists $i,i^{\prime}=1,\cdots,2N$ such that
@@
[…]
@@ Appendix B Numeric study of the 1-loop effect for $d>9$
At leading order of $y$, they both scale as $-2\log y$. However, this approximation breaks down quickly because as $n$ increases $ny/\gamma$ is no longer a small number.
@@
Figure 16: (a) Numeric plot of $-2\log Z_{h\neq 2}(\beta)$ for $d=10$ to $d=28$. (b) The fit parameters $a_{1}$ (blue) and $a_{2}$ (yellow) for $d=10$ to $d=28$. (c) Numeric plot of $\overline{-2\log Z_{h\neq 2}(\beta)}$, which has constant and linear fit parts subtracted, for $d=10$ to $d=28$. Note that the horizontal axis is in log scale.

### v4

@@ Appendix A Overlap of a generic pair of $X^{a}_{I}$ and $X^{b}_{I^{\prime}}$
For each pair $a\neq b$, we can first redefine $\psi_{i}$ appropriately such that the relative overlapping is equivalent to $X_{I}^{1}$ and $X_{I^{\prime}}^{2}$. To see this, consider $X_{I}^{a}$ and $X_{I^{\prime}}^{a+b}$ with $b>0$. We can first redefine $\psi_{2i}\rightarrow\psi_{2i-2(a-1)}$ to shift $X_{I}^{a}\rightarrow X_{I}^{1}$ and $X_{I}^{a+b}\rightarrow X_{I}^{1+b}$. We first assume $N$ is coprime with any number between $2$ and d1d-1. Then, we redefine $\psi_{2i-1}\rightarrow\psi_{2bi-(2b-1)}$ and $\psi_{2i}\rightarrow\psi_{2bi-(2b-2)}$. Since $b$ takes values between $1$ and d1d-1, by our coprime assumption this redefinition is a permutation with one cycle. Since random overlapping does not care about the explicit numbering of Majorana fermions, this problem is equivalent to the overlapping between $X_{I}^{1}$ and $X_{I^{\prime}}^{2}$.
For a generic case where $N$ is not coprime with a number between $2$ and d1d-1. The second redefinition does not lead to the overlapping problem between $X_{I}^{1}$ and $X_{I^{\prime}}^{2}$ but $b$ copies of overlapping problem with $N\rightarrow N/b$. To be precise, assume $N/b\in\mathbb{Z}$ and there exists $i,i^{\prime}=1,\cdots,2N$ such that
@@
where in the last step we used large $N$ approximation $\sum_{p_{i}}P(\{p_{i}\})\rightarrow\int dp_{i}\delta(p_{i}-p/b)$. This shows that the overlapping probability is universal in large $N$ limit. Nevertheless, if $d$ is very large and carefully chosen such that $b$ scales with $N$ or we define $\mathcal{X}_{i}^{a}$ alternatively with order $N$ numbers of relative cycles between two different $a$’s, the delta function approximation for $P(\{p_{i}\})$ does not hold and we have to consider the correction to this probability distribution.
@@ Appendix B Numeric study of the 1-loop effect for $d>9$
At leading order of $y$, they both scale as $2\log y$. However, this approximation breaks down quickly because as $n$ increases $ny/\gamma$ is no longer a small number.
@@
Figure 16: (a) Numeric plot of $2\log Z_{h\neq 2}(\beta)$ for $d=10$ to $d=28$. (b) The fit parameters $a_{1}$ (blue) and $a_{2}$ (yellow) for $d=10$ to $d=28$. (c) Numeric plot of $\overline{-2\log Z_{h\neq 2}(\beta)}$, which has constant and linear fit parts subtracted, for $d=10$ to $d=28$. Note that the horizontal axis is in log scale.